OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5).

OpenStudy (anonymous):

m= -1 , x1 = -2 , y1 =5 Substitute all the variables into the formula y-y1=m(x-x1) y-5 = -1(x-(-2)), expand it and rearrange y=-x+7